OpenStudy (anonymous):

What is the variance of the following data? If necessary, round your answer to two decimal places.12, 13, 8, 13, 9, 11, 12, 9, 11

Directrix (directrix):

What is the formula for variance of a data set? Are you allowed to use technology to get the answer?

OpenStudy (anonymous):

yes

Directrix (directrix):

If you are allowed to use technology, then go to the online variance calculator at the following link: http://easycalculation.com/statistics/standard-deviation.php Enter the data with no spaces as: 12,13,8,13,9,11,12,9,11 Please post what you get and we can compare answers.

Directrix (directrix):

@mariya18 What answer did you get?

OpenStudy (anonymous):

1/8=0.125 1/4=0.25 1/2=0.5 Explain how could you get from one row to another